# Infrared fixed point in the massless twelve-flavor SU(3) gauge-fermion system

### 28 Feb 2024 v1

@@ V Conclusions
We have calculated the non-perturbative $\beta$ function of the SU(3) gauge-fermion system with twelve massless fundamental fermions using a Pauli-Villars improved lattice action. We find strong evidence for an infrared fixed point at $g^{2}_{\mathrm{GF}\star}=6.69(68)$ from our gradient-flow-based renormalization scheme. Our study utilizes a wide range of couplings and volumes. In particular, we can reach renormalized gauge coupling values well above the predicted IRFP without the interference of a bulk phase transition. We include systematic effects from the infinite volume extrapolation directly into our analysis using Bayesian model averaging. Our data exhibits cutoff effects that are consistent with the leading $\mathcal{O}(a^{2}/t)$ form over a wide range $t/a^{2}$. The consistency between the W and C operators further supports the leading order scaling behavior. We believe the improved scaling is due to the additional PV bosons that reduce cutoff effects. In contrast, we found significantly larger cutoff effects when we re-analyzed the data that were generated without PV fields and used in Ref. Hasenfratz and Schaich (2018). Our data is publicly available at Ref. Peterson and Hasenfratz (2024).
Overall, the systematics of our continuum extrapolation are well controlled. Based on the continuum prediction for $\beta\big{(}g^{2}_{\mathrm{GF}}\big{)}$ in $g^{2}_{\mathrm{GF}}$, we estimate the leading irrelevant critical exponent $\gamma_{g}^{\star}$. This estimate includes systematic errors from various choices in our analysis. The value $\gamma_{g}^{\star}=0.204(36)$ agrees with Refs. Ryttov and Shrock (2016b); Di Pietro and Serone (2020) at the $1\sigma$ level and Ref. Hasenfratz and Schaich (2018) at the $2\sigma$ level.

### 13 Mar 2024 v2

@@ V Conclusions
We have calculated the non-perturbative $\beta$ function of the SU(3) gauge-fermion system with twelve massless fundamental fermions using a Pauli-Villars improved lattice action. We find strong evidence for an infrared fixed point at $g^{2}_{\mathrm{GF}\star}=6.60(62)$ from our gradient-flow-based renormalization scheme. Our study utilizes a wide range of couplings and volumes. In particular, we can reach renormalized gauge coupling values well above the predicted IRFP without the interference of a bulk phase transition. We include systematic effects from the infinite volume extrapolation directly into our analysis using Bayesian model averaging. Our data exhibits cutoff effects that are consistent with the leading $\mathcal{O}(a^{2}/t)$ form over a wide range $t/a^{2}$. The consistency between the W and C operators further supports the leading order scaling behavior. We believe the improved scaling is due to the additional PV bosons that reduce cutoff effects. In contrast, we found significantly larger cutoff effects when we re-analyzed the data that were generated without PV fields and used in Ref. Hasenfratz and Schaich (2018). Our data is publicly available at Ref. Peterson and Hasenfratz (2024).
Overall, the systematics of our continuum extrapolation are well controlled. Based on the continuum prediction for $\beta\big{(}g^{2}_{\mathrm{GF}}\big{)}$ in $g^{2}_{\mathrm{GF}}$, we estimate the leading irrelevant critical exponent $\gamma_{g}^{\star}=0.199(32)$. This estimate includes conservative systematic errors from various choices in our analysis. Our result for $\gamma_{g}^{\star}$ agrees with Refs. Ryttov and Shrock (2017); Di Pietro and Serone (2020) at the $1\sigma$ level and Ref. Hasenfratz and Schaich (2018) at the $2\sigma$ level.